A two hundred forty-four centimeter board is divided into two pieces. The longer piece is fourteen centimeters longer than four times the shorter piece. Find the length on each piece.

Answer by checkley79(3341)

X+Y=244
X=4Y+14
4Y+14+Y=244
5Y=244-14
5Y=230
Y=230/5
Y=46 cm IS THE SHORTER PIECE.
X=4*46+14
X=184+14
X=198 cm. FOR THE LONGER PIECE.
PROOF:
46+198=244
244=244